Suppose the risk-free rate is 2.62% and an analyst assumes a market risk premium of 6.31%. Firm A just paid a dividend of $1.48 per share. The analyst estimates the β of Firm A to be 1.27 and estimates the dividend growth rate to be 4.77% forever. Firm A has 263.00 million shares outstanding. Firm B just paid a dividend of $1.52 per share. The analyst estimates the β of Firm B to be 0.74 and believes that dividends will grow at 2.70% forever. Firm B has 195.00 million shares outstanding. What is the value of Firm B?

Round to 2 decimal places.
Answer #1

Using CAPM Model,

Required Rate of Stock B = 0.0262 + 0.74(0.0631)

Required Rate of Stock B = 7.29%

Using Constant Growth Model,

Stock Price = 1.52(1.027)/(0.0729 - 0.027)

Stock Price = $34.01

Value of Firm B = 195,000,000(34.01)

Value of Firm B = $6,631,869,281.10